Consider Young's experiment where two narrow slits are 0.40 mm apart from center to center. The irradiance at each slit is 0.25 W/m^2. Fringes separated by 3.40 mm appear on a screen 2.00 mm away. Write an expression for the irradiance at the screen as a function of x, the displacement from the central axis.
1
Expert's answer
2020-05-14T09:04:31-0400

As per the given question,

distance between the two slit (d)=0.4mm

Radiations "(I)=0.25W\/m^2"

The separation between the fringes y=3.40mm

screen distance D=2mm

"y=\dfrac{\lambda D}{d}"

"\Rightarrow \lambda = \dfrac{yd}{D}"

"\Rightarrow \lambda = \dfrac{3.40\times 0.4}{2}=0.68mm"

"\phi = \dfrac{2\pi}{\lambda}x=\dfrac{2\pi}{0.68}x"

"I=I_o \cos^2\phi=0.25 \cos(\dfrac{2\pi}{0.68})x"